OpenStudy (anonymous):

*help requested* *fan and medal awarded* Which relation does not represent a function? (-7, 3), (-3, 1), (1, -4), (5, -6) (-4, 2), (1, 5), (-4, 0), (5, 6) (2, 0), (4, 8), (6, 5), (10, 5) (-3, 6), (-1, 1), (0, 4), (3, 7) What is the slope and y-intercept of the graph of 2x + 3y = 18? slope = -2; y-intercept = (0, 18) slope = 2 ; y-intercept = (0, 6) - -- 3 slope = 3 ; y-intercept = (0, 2) - -- 2 slope = 6; y-intercept = (0, 2) - -- 3

OpenStudy (yrelhan4):

slope is -(coefficient of x)/(coefficient of y) and to find the y intercept put x=0 and solve..

OpenStudy (yrelhan4):

B becasue each number represents a domain and range number in a physical plain (-4,2) has a Domain of -4 and a range of 2 by laws of math, a real function CAN NOT HAVE THE SAME DOMAIN meaning that if any of the first numbers are the same in the set, it CAN NOT BE A FUNCTION so in B you have (-4, 2), (1, 5), (-4, 0), (5, 6) since (-4,2) and (-4,0) both share a -4 as their domain this set can not be a function

OpenStudy (anonymous):

math is gibrish to me. can u explain it in a simpler way ?

OpenStudy (anonymous):

For a simpler way to determine if points form a function you plot the points on a coordinate plane (graph the points). Then any time you can draw a lines up and down and if any line goes through more than one point you do not have a function.

OpenStudy (anonymous):

oh okay. but i think the first one is c and the second is b. am i right ?

OpenStudy (anonymous):

@novagirl114 @yrelhan4

OpenStudy (yrelhan4):

both B..

OpenStudy (anonymous):

explain please

OpenStudy (yrelhan4):

well i just explained.. slope is -(coefficient of x)/(coefficient of y) = -2/3 and for y intercept put x=0 and you'll get y=6

OpenStudy (yrelhan4):

and if any of the first numbers are the same in the set, it CAN NOT BE A FUNCTION.